OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2? A. y = 2x – 1 B. y = 2x + 1 C. y = –2x – 1 D. y = –2x + 1

OpenStudy (anonymous):

Y-Y1=m(x-X1) sub each term with its value so you get y-1=2(x-1) y=2x-2+1 y=2x-1 matches with A

OpenStudy (anonymous):

equation should be (y-1)=2(x-1) so y-1=2x-2--> y=2x-1...this is called slope intercept form